OpenStudy (anonymous):

A WAVE TRAVELS 10m IN 2s . what is the wave velocity ??

OpenStudy (hwyl):

velocity can be measured by the distance it takes per unit time generally \(m~s^{-1} \)

OpenStudy (hwyl):

waves' velocity is measured in "frequency" or Hertz (Hz)

OpenStudy (hwyl):

usually measured in per seconds :) so if you have 20 m in 2 seconds, all you have to do is simplify it so that denominator is 1 instead of 2

OpenStudy (legitemal):

The velocity formula is D/T.

OpenStudy (hwyl):

correct @Legitemal but the distance is measured in wavelength

OpenStudy (shamim):

Can u tell me the distance travelled by 1 sec

OpenStudy (anonymous):

just apply simple formula.."wave" is written to confuse u...wait which grade are u in? anyways it should be v=d/t ie v=10/2 = 5m/s